# Dataset Condensation with Color Compensation

### v2

@@ Abstract
Dataset condensation always faces a constitutive trade-off: balancing performance and fidelity under extreme compression. Existing methods struggle with two bottlenecks: image-level selection methods (Coreset Selection, Dataset Quantization) suffer from inefficiency condensation, while pixel-level optimization (Dataset Distillation) introduces semantic distortion due to over-parameterization. With empirical observations, we find that a critical problem in dataset condensation is the oversight of color’s dual role as an information carrier and a basic semantic representation unit. We argue that improving the colorfulness of condensed images is beneficial for representation learning. Motivated by this, we propose DC3: a Dataset Condensation framework with Color Compensation. After a calibrated selection strategy, DC3 utilizes the latent diffusion model to enhance the color diversity of an image rather than creating a brand-new one. Extensive experiments demonstrate the superior performance and generalization of DC3 that outperforms SOTA methods across multiple benchmarks. To the best of our knowledge, besides focusing on downstream tasks, DC3 is the first research to fine-tune pre-trained diffusion models with condensed datasets. The FID results prove that training networks with our high-quality datasets is feasible without model collapse or other degradation issues. Code and generated data are available at https://github.com/528why/Dataset-Condensation-with-Color-Compensation.
@@ 1 Introduction
In data-centric artificial intelligence, the dataset serves as a vehicle of knowledge whose information density and representation capacity determine the cognitive boundaries of a model. Information theory reveals a fundamental contradiction: There exists a coexistence of sparse task-relevant information and redundant task-irrelevant noise within raw data (Shannon, 1948; Cover, 1999; MacKay, 2003). Inspired by this contradiction, researchers propose dataset condensation, which focuses on creating a small dataset from the original large-scale data to reduce training costs while preserving comparable performance (Lei & Tao, 2023a; Yu et al., 2023a). The condensed dataset not only saves storage space but also accelerates I/O speed.
@@
As depicted in figs.˜2(a), 2(b) and 2(c), the RGB channels of the distilled datasets follow different distributions compared to the original datasets (especially for the matching surrogate objective methods (Zhao & Bilen, 2023; 2021a; Sajedi et al., 2023). Their KDE curves all tend to have a uniform distribution. Such Color Homogenization phenomenon could hinder the model in understanding the data representations, thus degrading its generalization. Unlike DD, quantized datasets are derived through a quantization process from the original data, notably preserving precise color distribution and semantic integrity, whereas lacking the information compression capability inherent to DD. Motivated by these observations, we seek to study: “Can we avoid Color Homogenization, but attain an informative dataset by directly using the image-level selection?
@@
Recent advancements in generative modeling have established diffusion models as a dominant approach for text-to-image and image-to-image generation. Building upon these successes, pioneering works have adapted diffusion models to dataset condensation, leveraging their high-fidelity generation capability to preserve essential characteristics within compressed datasets (Su et al., 2024; Cazenavette et al., 2023; Gu et al., 2024). However, diffusion models exhibit limited robustness under sparse or ambiguous textual conditions, generating images with semantic inconsistency or hallucination that deviate from real-world distributions.
@@
  • We propose DC3 that utilizes the compression ability as long as the generalization from different dataset condensation methods. It is a universal method that is adaptive to datasets of various scales and resolutions.
@@ 2.1 Dataset Distillation
Dataset Distillation (DD) (Wang et al., 2018; Li et al., 2022; Lei & Tao, 2023b; Yu et al., 2023b) methods are categorized into matching-based and generation-based approaches. Matching-based methods can also be categorized into three paradigms: performance matching, parameter matching, and distribution matching. Performance matching focuses on optimizing the synthetic dataset to ensure that models trained on it perform similarly to those trained on the original dataset (Nguyen et al., 2021a; b). Parameter matching, which includes single-step (e.g., gradient matching (Zhao & Bilen, 2021b; a)) and multi-step (e.g., trajectory matching (Cazenavette et al., 2022; Guo et al., 2024)) approaches, aims to align the model parameters or their learning trajectories during training on synthetic and real data. Distribution matching emphasizes narrowing the distribution gap by ensuring that synthetic data approximates the feature distribution of the original data (Zhao & Bilen, 2023; Deng et al., 2024).
@@ 2.2 Dataset Quantization
Traditional DD methods often rely on specific network architectures, resulting in degraded performance when the distilled datasets are transferred to other architectures (Wang et al., 2018; Zhao & Bilen, 2021b; Cazenavette et al., 2022). To address this limitation, Dataset Quantization (DQ) (Zhou et al., 2023) introduces a novel framework capable of compressing large-scale datasets and generating compact datasets suitable for training any neural network architecture without sacrificing performance. DQ achieves this by recursively partitioning the dataset into non-overlapping data bins and optimizing sample selection based on submodular gains to maximize dataset diversity. This approach not only enhances dataset diversity but also ensures the broad applicability of the generated subsets across varying network architectures. Zhao et al. (2024) further advanced this paradigm by proposing DQAS, an adaptive sampling strategy integrating active learning to refine class-imbalanced sample selection. Additionally, this work introduces a new pipeline utilizing feature spaces from advanced quantization stages to generate higher-precision data bins. This methodology addresses the limitations of traditional DQ approaches that assume uniform class distributions, enabling more efficient knowledge transfer across heterogeneous network architectures. Similarly, Li et al. (2025) proposes ADQ to improve naive DQ by adaptively sampling data based on the representativeness, diversity, and importance scores of the generated bins.
@@ 3.1 Image Processing with Latent Diffusion
where $\mathcal{E}(x)$ and $\tau_{\theta}(y)$ represent the latent embeddings of the target image $x$ and conditional prompt $y$, encoded by a pretrained autoencoder $\mathcal{E}$ and CLIP $\tau_{\theta}$. At each timestep, noise $\epsilon$ is added to the latent $z_{t}$, and the denoising network $\epsilon_{\theta}$ learns to predict the noise conditioned on $y$.
@@ 3.2 Color Compensation for Dataset Condensation
A hue is the dominant color family of a particular color. In DC3, the hue prompts are categorized into cool (e.g., rainy, snowy) and warm (e.g., sunny, daylight), i.e., except for the neutral colors (e.g., black, white, and gray). We select two hues from each group: $(\mathcal{P}_{\texttt{cool}},\mathcal{P}_{\texttt{warm}})$, and implement the color compensation to synthesize $(I_{\texttt{cool}},I_{\texttt{warm}})$ according to eq.˜2. Please refer to appendix˜C for a detailed guideline on prompt selection. Furthermore, we crop them in half and then stitch them to create a more informative image.
As illustrated in fig.˜4, color compensation effectively enhances dataset colorfulness by expanding the color gamut representation without introducing artificial artifacts, thereby contributing to model training on downstream tasks. Compared to RDED (Sun et al., 2024), another crop-and-stitch approach that composites images through disparate image regions, our method synthesizes multiple color-compensated variants of identical images. This strategy preserves structural consistency while promoting generalized representations across diverse visual domains such as resolution and color distribution.
@@ 3.3 Quantization Analysis
Figure 5: T-SNE visualizations of intra-class clusters on ImageNet. Submodular sampling selects the representative samples to preserve maximum semantic integrity and feature diversity.
DQ is an emerging method for dataset condensation based on image-level selection. The idea of DQ is to partition data into different bins iteratively based on submodular gain, which was first proposed in GraphCut (Iyer et al., 2021) for coreset selection. The samples are selected from the bins randomly, whose distribution is aligned with the original datasets, i.e., it does not have the color homogenization problem. Nevertheless, a fundamental limitation of this approach lies in its inherently monotonic decay of submodular gains across increasing bin indices, which inevitably causes random sampling to discard critical samples essential for model generalization. To reveal the relationship between submodular gain and sample distribution, we visualize the samples in ImageNette according to the values of their submodular gains with the help of t-SNE (Van der Maaten & Hinton, 2008) plots (the mid-row of fig.˜6). Also, we provide additional visualizations for the first, second, and last samples selected by the submodular gain criterion (the top row and bottom row of fig.˜6).
Figure 6: Visualizations of the relationship between submodular gain and sample distribution. The following situations occur in traditional DQ methods: (a) inconsequential but compact samples may be selected because they are in different bins (bottom-left), and (b) essential and diverse samples may be discarded because they are in the same bins (top-right).
Due to DQ performing random sampling within a certain bin, the following situations exist: (a) Samples with high submodular gains may be excluded despite their distance from the optimal sample (top-right in fig.˜6). (b) Samples with low submodular gains may be selected even when proximate to the optimal sample (bottom-left in fig.˜6). Based on the visualization, we argue that the submodular gain of a sample is independent of its position in the feature distribution. In other words, the sampling diversity does not necessarily benefit submodular gain.
To avoid selecting samples with small gains, we propose clustering-based DQ, where the submodular gain is then used as an image sampler instead of a bin generator. Specifically, the clustering algorithm divides the dataset into multiple clusters, which are naturally regarded as data bins ($C_{j}$). Each sample $x$ is iteratively assigned to the closest bins:
@@
As shown in fig.˜5, data bins, also known as clusters, are dispersed throughout the feature space as expected. The clustering algorithm performs a coarse partitioning of samples based on feature-space similarities. Within each bin, we leverage submodular sampling to identify the representative samples with high submodular gains. In this way, the condensed dataset preserves maximum semantic and feature diversity.
@@ 4.1 Experimental Settings
We perform systematic evaluations across varying-scale benchmarks to assess generalization boundaries. For large-scale datasets, we include ImageNet-1K (Deng et al., 2009) ($224\times 224$) and its subsets, such as Tiny-ImageNet ($64\times 64$). Small-scale low-resolution ($32\times 32$) analysis utilizes CIFAR-10/100 (Krizhevsky et al., 2009). To quantify task difficulty sensitivity, we benchmark ImageNette and ImageWoof, the subsets of ImageNet with 10 classes. Note that ImageWoof poses greater discrimination challenges due to more inter-class similarity.
@@ 4.2 Comparison with SOTA Methods
While our method draws inspiration from DQ, we exclude direct comparisons due to differing evaluation priorities. Specifically, DQ emphasizes generalization across large IPC keep ratios (e.g., 10%, 20%, 30%), whereas our focus lies in optimizing compression performance through color compensation at low IPC regimes. Consequently, we benchmark against DD such as the matching-based TESLA (Cui et al., 2023), SRe2(Yin et al., 2023), DataDAM, RDED (Sun et al., 2024), and generation-based Minimax (Gu et al., 2024), D4(Su et al., 2024), and IGD (Chen et al., 2025), aligning with the objective of DC that maximizes dataset utility under extreme compression. Importantly, our DC3 inherits all the advantages of DQ while achieving superior compression efficiency.
@@ 4.3 Cross-architecture Generalization
To support the claim that pixel optimization methods, which cause “color homogenization”, impair model generalization, a direct comparison is made between DC3 and various matching-based DD methods (such as DM, DSA) on “Seen” and “Unseen” network architectures. The results in table˜4 indicate that the aforementioned methods experience a sharp accuracy decrease when evaluated on architectures not previously seen. Such significant performance degradation provides strong empirical support for the hypothesis that “color homogenization” hinders model learning of transferable data representations. In contrast, DC3 demonstrates the smallest performance degradation on the unseen architecture, at only 9.8%, and its absolute performance has far exceeded that of other methods.
Eval. Model DM DSA MTT TESLA DC3
Seen $63.0$ $60.6$ $65.3$ $66.4$ 89.8
Unseen $52.9_{{\color[rgb]{0,1,1}\downarrow 10.1}}$ $47.8_{{\color[rgb]{0,1,1}\downarrow 12.8}}$ $34.6_{{\color[rgb]{0,1,1}\downarrow 30.7}}$ $34.8_{{\color[rgb]{0,1,1}\downarrow 31.6}}$ $\textbf{80.0}_{{\color[rgb]{0,1,1}\downarrow 9.8}}$
Table 4: Different DD methods show performance degradation on the “unseen” architecture. We choose the worst result of DC3 on Swin-V2-T under IPC-50, but it still performs the best among methods.
@@
ResNet-18 10 84.8 80.8 79.0 81.8 84.2 60.4
50 89.8 89.0 89.2 89.6 90.2 80.0
@@ 4.4 Ablation Study
CIFAR10 10 50.3 54.5 57.8
50 73.6 76.0 80.9
CIFAR100 10 51.8 53.5 57.4
50 58.3 61.2 64.2
@@
Specifically, through the red-green (rg) and yellow-blue (yb) difference of an image, the colorfulness metric can be defined as:
@@
As depicted in fig.˜7, the application of color compensation techniques to submodular-sampled datasets demonstrates significant theoretical advantages over uncompensated approaches. The compensation strategy systematically addresses the inherent color distribution imbalances in compressed datasets by enhancing the color of underrepresented regions while maintaining visual naturalness. The cool hue specifically targets cyan-blue-magenta deficiencies, while warm compensation addresses yellow-orange-red limitations. Thus, the mixed compensation provides a comprehensive spectral coverage. Appendix˜D discusses the different mix methods in detail.
Compensation v.s. augmentation. To comprehensively evaluate the dual contributions of our proposed DC3 framework in color compensation and coreset selection, a detailed ablation study is conducted with results listed in table˜7. Since one of the primary contributions of DC3 is addressing the issue of color homogenization in existing methods, table˜7 directly compares DC3 against other color augmentation techniques such as Color Jitter, HSV Shift, Balance Tuning, and AutoPalette Yuan et al. (2024). The experimental results demonstrate that DC3 significantly outperforms these alternative color solutions on the CIFAR-10 and CIFAR-100 datasets, thus confirming the unique advantage of the diffusion model-based approach in enhancing data color diversity. We named this strategy compensation instead of augmentation to distinguish it.
@@
CIFAR10 73.5 74.2 74.6 79.4 80.9
CIFAR100 53.2 53.7 54.4 53.3 64.2
Table 7: Comparison with augmentation methods under IPC-50.
Submodular sampling. Table˜8 demonstrates that clustering-based DQ with submodular sampling performs better across both CIFAR-10 and ImageNette under varying IPCs. Specifically, we achieve notable improvements of 3.1% and 2.9% at IPC-10 on CIFAR-10 and ImageNette, respectively. These consistent performance enhancements across different compression ratios validate the effectiveness of the proposed methodological refinements.
@@
IPC = 10 IPC = 50
DQ (w/o clustering) 46.3 72.4 69.4 76.8
DC3 (w/ clustering) $\textbf{49.4}^{{\color[rgb]{1,.5,0}\uparrow 3.1}}$ $\textbf{75.3}^{{\color[rgb]{1,.5,0}\uparrow 2.9}}$ $\textbf{71.4}^{{\color[rgb]{1,.5,0}\uparrow 2.0}}$ $\textbf{79.6}^{{\color[rgb]{1,.5,0}\uparrow 2.8}}$
Table 8: Ablation study on sampling strategies. DQ randomly selects samples within data bins, while DC3 utilizes submodular gain as the sampler. $\dagger$: Considering the orthogonality of experimental design, DC3 performs evaluation directly after sampling in this experiment.
@@
- - 38.5 55.8
- $49.4^{{\color[rgb]{1,.5,0}\uparrow 10.9}}$ $75.3^{{\color[rgb]{1,.5,0}\uparrow 19.5}}$
- $53.6^{{\color[rgb]{1,.5,0}\uparrow 15.1}}$ $77.9^{{\color[rgb]{1,.5,0}\uparrow 22.1}}$
$\textbf{57.8}^{{\color[rgb]{1,.5,0}\uparrow 19.3}}$ $\textbf{84.8}^{{\color[rgb]{1,.5,0}\uparrow 29.0}}$

### v3

@@ Abstract
Dataset condensation always faces a constitutive trade-off: balancing performance and fidelity under extreme compression. Existing methods struggle with two bottlenecks: image-level selection methods (Coreset Selection, Dataset Quantization) suffer from inefficiency in condensation, while pixel-level optimization (Dataset Distillation) introduces semantic distortion due to over-parameterization. With empirical observations, we find that a critical problem in dataset condensation is the oversight of color’s dual role as an information carrier and a basic semantic representation unit. We argue that improving the colorfulness of condensed images is beneficial for representation learning. Motivated by this, we propose DC3: a Dataset Condensation framework with Color Compensation. After a calibrated selection strategy, DC3 utilizes the latent diffusion model to enhance the color diversity of an image rather than creating a brand-new one. Extensive experiments demonstrate the superior performance and generalization of DC3 that outperforms SOTA methods across multiple benchmarks. To the best of our knowledge, besides focusing on downstream tasks, DC3 is the first research to fine-tune pre-trained diffusion models with condensed datasets. The Frechet Inception Distance (FID) and Inception Score (IS) results prove that training networks with our high-quality datasets is feasible without model collapse or other degradation issues. Code and generated data are available at https://github.com/528why/Dataset-Condensation-with-Color-Compensation.
@@ 1 Introduction
In data-centric artificial intelligence, the dataset serves as a vehicle of knowledge whose information density and representation capacity determine the cognitive boundaries of a model. Information theory reveals a fundamental contradiction: There exists a coexistence of sparse task-relevant information and redundant task-irrelevant noise within raw data (Shannon, 1948; Cover, 1999; MacKay, 2003). Inspired by this contradiction, researchers propose dataset condensation, which focuses on creating a small dataset from the original large-scale data to reduce training costs while preserving comparable performance (Geng et al., 2023; Lei & Tao, 2023a; Yu et al., 2023a). The condensed dataset not only saves storage space but also accelerates I/O speed.
@@
As depicted in figs.˜2(a), 2(b) and 2(c), the RGB channels of the distilled datasets follow different distributions compared to the original datasets (especially for the matching surrogate objective methods (Zhao & Bilen, 2023; 2021a; Sajedi et al., 2023). Their Kernel Density Estimation (KDE) curves all tend to have a uniform distribution. Such Color Homogenization phenomenon could hinder the model in understanding the data representations, thus degrading its generalization. Unlike DD, quantized datasets are derived through a quantization process from the original data, notably preserving precise color distribution and semantic integrity, whereas lacking the information compression capability inherent to DD. Motivated by these observations, we seek to study: “Can we avoid Color Homogenization, but attain an informative dataset by directly using the image-level selection?
@@
Recent advancements in generative modeling have established diffusion models as a dominant approach for text-to-image and image-to-image generation. Building upon these successes, pioneering works have adapted diffusion models to dataset condensation, leveraging their high-fidelity generation capability to preserve essential characteristics within compressed datasets (Su et al., 2024; Cazenavette et al., 2023; Gu et al., 2024; Sajedi et al., 2024). However, diffusion models exhibit limited robustness under sparse or ambiguous textual conditions, generating images with semantic inconsistency or hallucination that deviate from real-world distributions.
@@
  • According to the dataset condensation triangle, we propose DC3 that utilizes the compression ability of dataset distillation along with the generalization ability of dataset quantization. It is a universal method that is adaptive to datasets of various scales and resolutions.
@@ 2.1 Dataset Distillation
Dataset Distillation (DD) (Wang et al., 2018; Li et al., 2022; Lei & Tao, 2023b; Yu et al., 2023b) methods are categorized into matching-based and generation-based approaches. Matching-based methods can also be categorized into three paradigms: performance matching, parameter matching, and distribution matching. Performance matching focuses on optimizing the synthetic dataset to ensure that models trained on it perform similarly to those trained on the original dataset (Nguyen et al., 2021a; c). Parameter matching, which includes single-step (e.g., gradient matching (Zhao & Bilen, 2021b; a)) and multi-step (e.g., trajectory matching (Cazenavette et al., 2022; Guo et al., 2024)) approaches, aims to align the model parameters or their learning trajectories during training on synthetic and real data. Distribution matching emphasizes narrowing the distribution gap by ensuring that synthetic data approximates the feature distribution of the original data (Zhao & Bilen, 2023; Deng et al., 2024).
@@ 2.2 Dataset Quantization
Traditional DD methods often rely on specific network architectures, resulting in degraded performance when the distilled datasets are transferred to other architectures (Wang et al., 2018; Zhao & Bilen, 2021b; Cazenavette et al., 2022). To address this limitation, Dataset Quantization (DQ) (Zhou et al., 2023) introduces a novel framework capable of compressing large-scale datasets and generating compact datasets suitable for training any neural network architecture without sacrificing performance. DQ achieves this by recursively partitioning the dataset into non-overlapping data bins and optimizing sample selection based on submodular gains to maximize dataset diversity. This approach not only enhances dataset diversity but also ensures the broad applicability of the generated subsets across varying network architectures. In a parallel effort to enhance both the diversity and realism of distilled data, Sun et al. (2024) introduces RDED, an efficient optimization-free paradigm that synthesizes new images by selecting and stitching patches from real data. Zhao et al. (2024) further advanced this paradigm by proposing DQAS, an adaptive sampling strategy integrating active learning to refine class-imbalanced sample selection. This methodology addresses the limitations of traditional DQ approaches that assume uniform class distributions, enabling more efficient knowledge transfer across heterogeneous network architectures. Similarly, Li et al. (2025) proposes ADQ to improve naive DQ by adaptively sampling data based on the representativeness, diversity, and importance scores of the generated bins.
@@ 3.1 Image Processing with Latent Diffusion
where $\mathcal{E}(x)$ and $\tau_{\theta}(y)$ represent the latent embeddings of the target image $x$ and conditional prompt $y$, encoded by a pretrained autoencoder $\mathcal{E}$ and CLIP $\tau_{\theta}$ (Radford et al., 2021). At each timestep, noise $\epsilon$ is added to the latent $z_{t}$, and the denoising network $\epsilon_{\theta}$ learns to predict the noise conditioned on $y$.
@@ 3.2 Color Compensation for Dataset Condensation
A hue is the dominant color family of a particular color. In DC3, the hue prompts are categorized into cool (e.g., rainy, snowy) and warm (e.g., sunny, daylight), except for the neutral hues (e.g., black, white, and gray). We select two hues from each group: $(\mathcal{P}_{\texttt{cool}},\mathcal{P}_{\texttt{warm}})$, and implement the color compensation to synthesize $(I_{\texttt{cool}},I_{\texttt{warm}})$ according to eq.˜2. Please refer to appendix˜C for a detailed guideline on prompt selection. Furthermore, we crop them in half and then stitch them to create a more informative image.
As illustrated in fig.˜4, color compensation effectively enhances dataset colorfulness by expanding the color gamut representation without introducing artificial artifacts, thereby contributing to model training on downstream tasks. Compared to RDED (Sun et al., 2024), another crop-and-stitch approach that composites images through disparate image regions, our method synthesizes multiple color-compensated variants of identical images. This strategy preserves structural consistency while promoting generalized representations across diverse visual domains, such as resolution and color distribution.
@@ 3.3 Quantization Analysis
DQ is an emerging method for dataset condensation based on image-level selection. The idea of DQ is to partition data into different bins iteratively based on submodular gain, which was first proposed in GraphCut (Iyer et al., 2021) for coreset selection. The samples are selected from the bins randomly, whose distribution is aligned with the original datasets, i.e., it does not have the color homogenization problem. Nevertheless, a fundamental limitation of this approach lies in its inherently monotonic decay of submodular gains across increasing bin indices, which inevitably causes random sampling to discard critical samples essential for model generalization. To reveal the relationship between submodular gain and sample distribution, we visualize the samples in ImageNette according to the values of their submodular gains with the help of t-SNE (Van der Maaten & Hinton, 2008) plots (the mid-row of fig.˜5). Also, we provide additional visualizations for the first, second, and last samples selected by the submodular gain criterion (the top row and bottom row of fig.˜5).
Figure 5: Visualizations of the relationship between submodular gain and sample distribution. The following situations occur in traditional DQ methods: (a) inconsequential but compact samples may be selected because they are in different bins (bottom-left), and (b) essential and diverse samples may be discarded because they are in the same bins (top-right). Each column represents a different category.
Due to DQ performing random sampling within a certain bin, the following situations exist: (a) Samples with high submodular gains may be excluded despite their distance from the optimal sample that has the highest submodular gain (top-right in fig.˜5). (b) Samples with low submodular gains may be selected even when proximate to the optimal sample (bottom-left in fig.˜5). Based on the visualization, we argue that the submodular gain of a sample is independent of its position in the feature distribution. In other words, the sampling diversity does not necessarily benefit submodular gain.
To avoid selecting samples with small gains, we propose clustering-based DQ, where the submodular gain is then used as an image sampler instead of a bin generator. Specifically, the clustering algorithm divides the dataset into multiple clusters, which are naturally regarded as data bins ($C_{j}$).
Figure 6: T-SNE visualizations of intra-class clusters on ImageNet. Submodular sampling selects the representative samples to preserve maximum feature diversity. Each column represents a different category.
Each sample $x$ is iteratively assigned to the closest bins:
@@
As shown in fig.˜6, data bins, also known as clusters, are dispersed throughout the feature space as expected. The clustering algorithm performs a coarse partitioning of samples based on feature-space similarities. Within each bin, we leverage submodular sampling to identify the representative samples with high submodular gains. In this way, the condensed dataset preserves maximum semantic and feature diversity.
@@ 4.1 Experimental Settings
We perform systematic evaluations across varying-scale benchmarks to assess generalization boundaries. For large-scale datasets, we include ImageNet-1K (Deng et al., 2009) ($224\times 224$) and its subsets, such as Tiny-ImageNet ($64\times 64$). Small-scale low-resolution ($32\times 32$) analysis utilizes CIFAR-10/100 (Krizhevsky et al., 2009). To quantify task difficulty sensitivity, we benchmark ImageNette and ImageWoof, the subsets of ImageNet with 10 classes. Note that ImageWoof poses greater discrimination challenges due to more inter-class similarity. We use Stable Diffusion-V1.5 and DiT-XL/2-256 as our foundation models.
@@ 4.2 Comparison with SOTA Methods
While our method draws inspiration from DQ, we exclude direct comparisons due to differing evaluation priorities. Specifically, DQ emphasizes generalization across large IPC keep ratios, whereas our focus lies in optimizing compression performance through color compensation at low IPC regimes. Consequently, we benchmark against DD such as the matching-based TESLA (Cui et al., 2023), SRe2(Yin et al., 2023), DataDAM (Sajedi et al., 2023), RDED (Sun et al., 2024), and generation-based Minimax (Gu et al., 2024), D4(Su et al., 2024), and IGD (Chen et al., 2025), aligning with the objective of DC that maximizes dataset utility under extreme compression. Importantly, our DC3 inherits all the advantages of DQ while achieving superior compression efficiency.
@@ 4.3 Cross-architecture Generalization
To support the claim that pixel optimization methods, which cause “color homogenization”, impair model generalization, a direct comparison is made between DC3 and various matching-based DD methods (such as KIP (Nguyen et al., 2021b), DSA) on “Seen” and “Unseen” network architectures. The results in table˜4 indicate that the aforementioned methods experience a sharp accuracy decrease when evaluated on architectures not previously seen. Such significant performance degradation provides strong empirical support for the hypothesis that “color homogenization” hinders model learning of transferable data representations. Among all methods, DC3 and DSA demonstrate the most robust generalization to the unseen architecture, exhibiting the slightest performance degradation. Crucially, DC3 absolute performance has far exceeded that of other methods.
Eval. Model KIP DSA MTT TESLA DC3
Seen $47.2$ $53.0$ $65.3$ $66.4$ $\textbf{84.8}_{\pm 1.2}$
Unseen $15.9_{{\color[rgb]{0,1,1}\downarrow 31.3}}$ $31.9_{{\color[rgb]{0,1,1}\downarrow 21.1}}$ $34.6_{{\color[rgb]{0,1,1}\downarrow 30.7}}$ $34.8_{{\color[rgb]{0,1,1}\downarrow 31.6}}$ $\textbf{60.4}_{{\pm 1.2}_{{\color[rgb]{0,1,1}\downarrow 24.4}}}$
Table 4: Different DD methods show performance degradation on the “unseen” architecture. Here, “seen” architectures refer to CNN-based models (e.g., ResNet-18 or ConvNet-D6), while “unseen” architectures are Transformer-based (e.g., Swin-V2-T or ViT-S). We choose the worst result of DC3 on Swin-V2-T under IPC-10, but it still performs the best among methods.
@@
ResNet-18 10 $\textbf{84.8}_{\pm 1.2}$ $80.8_{\pm 1.5}$ $79.0_{\pm 1.3}$ $81.8_{\pm 1.1}$ $\underline{84.2}_{\pm 1.4}$ $60.4_{\pm 1.2}$
50 $\underline{89.8}_{\pm 0.3}$ $89.0_{\pm 0.2}$ $89.2_{\pm 0.4}$ $89.6_{\pm 0.5}$ $\textbf{90.2}_{\pm 0.4}$ $80.0_{\pm 0.5}$
@@ 4.4 Ablation Study
CIFAR-10 10 $50.3_{\pm 0.2}$ $54.5_{\pm 0.2}$ $\textbf{57.8}_{\pm 0.8}$
50 $73.6_{\pm 0.1}$ $76.0_{\pm 0.2}$ $\textbf{80.9}_{\pm 0.5}$
CIFAR-100 10 $51.8_{\pm 0.1}$ $53.5_{\pm 0.2}$ $\textbf{57.4}_{\pm 0.2}$
50 $58.3_{\pm 0.1}$ $61.2_{\pm 0.1}$ $\textbf{64.2}_{\pm 0.5}$
@@
Specifically, through the red-green (rg) and yellow-blue (yb) difference of an image, the colorfulness can be defined as:
@@
As depicted in fig.˜7, the application of color compensation to submodular-sampled datasets demonstrates significant theoretical advantages over uncompensated approaches. The compensation strategy addresses the inherent color distribution imbalances in compressed datasets by enhancing the color of underrepresented regions while maintaining visual naturalness. The cool hue specifically targets cyan-blue-magenta deficiencies, while warm compensation addresses yellow-orange-red limitations. Thus, the mixed compensation provides a comprehensive spectral coverage. Appendix˜D discusses the different mix methods in detail.
Compensation v.s. augmentation. To comprehensively evaluate the dual contributions of our proposed DC3 framework in color compensation and coreset selection, a detailed ablation study is conducted with results listed in table˜7. Since one of the primary contributions of DC3 is addressing the issue of color homogenization in existing methods, table˜7 directly compares DC3 against other color augmentation techniques such as Color Jitter, HSV Shift, Balance Tuning, and AutoPalette (Yuan et al., 2024). The results demonstrate that DC3 significantly outperforms these alternative color solutions on the CIFAR-10 and CIFAR-100 datasets, thus confirming the unique advantage of the diffusion-based approach in enhancing color diversity. We named this strategy compensation instead of augmentation to distinguish it.
@@
CIFAR-10 $73.5_{\pm 1.8}$ $74.2_{\pm 1.1}$ $74.6_{\pm 1.4}$ $79.4_{\pm 2.2}$ $\textbf{80.9}_{\pm 0.5}$
CIFAR-100 $53.2_{\pm 1.6}$ $53.7_{\pm 1.4}$ $54.4_{\pm 1.7}$ $53.3_{\pm 1.3}$ $\textbf{64.2}_{\pm 0.8}$
Table 7: Comparison with augmentation methods under IPC-50. For fair comparison, we substitute Colour Compensation with augmentation methods, while the rest of the pipeline remains identical.
Submodular sampling. The upper section of table˜8 reveals that DC3 with clustering-based submodular sampling performs better across CIFAR-10 and ImageNette under varying IPCs. The lower section only changes the sampling strategy, while the rest of the components remain fixed, including color compensation.
@@
Method Color Compensation IPC = 10 IPC = 50
DQ 35.7±0.7 69.3±0.5 $50.5_{\pm 0.4}$ $74.2_{\pm 0.5}$
DC3 $\textbf{49.4}_{{\pm 0.7}^{{\color[rgb]{1,.5,0}\uparrow 13.7}}}$ $\textbf{75.3}_{{\pm 0.6}^{{\color[rgb]{1,.5,0}\uparrow 6.0}}}$ $\textbf{71.4}_{{\pm 0.5}^{{\color[rgb]{1,.5,0}\uparrow 20.9}}}$ $\textbf{79.6}_{{\pm 0.4}^{{\color[rgb]{1,.5,0}\uparrow 5.4}}}$
DQ $51.2_{\pm 0.6}$ $81.8_{\pm 1.5}$ $68.4_{\pm 0.8}$ $85.3_{\pm 0.6}$
DC3 $\textbf{57.8}_{{\pm 0.8}^{{\color[rgb]{1,.5,0}\uparrow 6.6}}}$ $\textbf{84.8}_{{\pm 1.1}^{{\color[rgb]{1,.5,0}\uparrow 3.0}}}$ $\textbf{80.9}_{{\pm 0.5}^{{\color[rgb]{1,.5,0}\uparrow 12.5}}}$ $\textbf{89.8}_{{\pm 0.2}^{{\color[rgb]{1,.5,0}\uparrow 4.5}}}$
Table 8: Ablation study on sampling strategies. DQ randomly selects samples within data bins, while DC3 utilizes submodular gain as the sampler.
@@
- - $38.5_{\pm 0.1}$ $55.8_{\pm 0.1}$
- $49.4_{{\pm 0.3}^{{\color[rgb]{1,.5,0}\uparrow 10.9}}}$ $75.3_{{\pm 0.4}^{{\color[rgb]{1,.5,0}\uparrow 19.5}}}$
- $53.6_{{\pm 0.6}^{{\color[rgb]{1,.5,0}\uparrow 15.1}}}$ $77.9_{{\pm 0.9}^{{\color[rgb]{1,.5,0}\uparrow 22.1}}}$
$\textbf{57.8}_{{\pm 0.8}^{{\color[rgb]{1,.5,0}\uparrow 19.3}}}$ $\textbf{84.8}_{{\pm 1.1}^{{\color[rgb]{1,.5,0}\uparrow 29.0}}}$